Calculus I, Fall 2013 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 05

1. Section 3.8: problems 7, 9, 37, 39, 51 (read example 6 first), 65, 77, 89, 91, 98. 93, 95.

2. Section 3.9: problems 21, 23, 25, 33, 35, 39, 55.

3. Start with domain and range for csc and csc^{−1}, derive the formula for the derivative of csc^{−1}.